### v2

# Revisiting the Reliability of Language Models in Instruction-Following

Abstract
Advanced LLMs have achieved near-ceiling instruction-following accuracy on benchmarks such as IFEval. However, these impressive scores do not necessarily translate to reliable services in real-world use, where users often vary their phrasing, contextual framing, and task formulations. In this paper, we study nuance-oriented reliability: whether models exhibit consistent competence across cousin prompts that convey analogous user intents but with subtle nuances. To quantify this, we introduce a new metric, reliable@k, and develop an automated pipeline that generates high-quality cousin prompts via data augmentation. Building upon this, we construct IFEval++ for systematic evaluation. Across 20 proprietary and 26 open-source LLMs, we find that current models exhibit substantial insufficiency in nuance-oriented reliability—their performance can drop by up to 61.8% with nuanced prompt modifications. What’s more, we characterize it and explore three potential improvement recipes. Our findings highlight nuance-oriented reliability as a crucial yet underexplored next step toward more dependable and trustworthy LLM behavior. Our code and benchmark are accessible: https://github.com/jianshuod/IFEval-pp.
Revisiting the Reliability of Language Models in Instruction-Following
Jianshuo Dong1, Yutong Zhang1, Yan Liu2, Zhenyu Zhong2, Tao Wei2, Chao Zhang1, Han Qiu1thanks: Corresponding author. 1Tsinghua University, China. 2Ant Group, China.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable abilities to follow natural language instructions (Ouyang et al., 2022). The instruction-following ability is foundational for enabling faithful user interaction (Zhou et al., 2023; Lior et al., 2025; Zhang et al., 2024b), reliable agent behavior (Zhang et al., 2025a), and reduced harmful outcomes (Ruan et al., 2024). Evaluating this ability is essential for building reliable, trustworthy AI systems. Consequently, numerous benchmarks have been proposed to measure instruction-following performance across diverse task types and constraint categories (see Appendix˜A for a survey of 36 benchmarks).
Figure 1: Evolution of OpenAI’s Models in Instruction-Following Abilities Across Generations.
As illustrated in Figure˜1, the instruction-following ability of LLMs has improved rapidly across generations. Recent LLMs achieve near-ceiling performance on widely adopted benchmarks like IFEval (Zhou et al., 2023), with GPT-5 reaching an accuracy as high as 95.9%. However, growing evidence reveals that model performance can be highly sensitive to prompt wording (Sclar et al., 2024; Mizrahi et al., 2024; Cao et al., 2024), and that overfitting to benchmark test cases is possible (Zhang et al., 2024a). Thus, questions remain about how well benchmark accuracy translates to real-world reliability.
To illustrate this concern, consider two LLMs that both achieve perfect benchmark accuracy. When faced with user prompts expressing similar intent but differing subtly in phrasing, contextual framing, or task instantiation, one model may generalize appropriately while the other fails. Current benchmarks, which primarily emphasize task and constraint diversity, cannot capture this crucial nuance-oriented dimension of reliability. This motivates our central question: Are current LLMs reliable when handling “cousin prompts” that convey similar user intents but vary in nuanced ways?
Figure 2: Examples of Cousin Prompts. The corresponding evaluation configurations are provided.
To operationalize this, we propose a new metric, reliable@k, which quantifies whether an LLM can simultaneously solve a set of cousin prompts. We further design an automated pipeline for generating high-quality cousin prompts using three data augmentation strategies: rephrasing, distractor addition, and task/constraint reconfiguration. This is coupled with a code-assisted validity checker. Building on IFEval (Zhou et al., 2023), we leverage the pipeline to construct an extended benchmark, IFEval++ ***We name it to show respect for the renowned IFEval benchmark, as we position IFEval++ as a successor of IFEval, similar to how C++ relates to C. , which consists of 541 test cases, each comprising 10 cousin prompts. This provides a testbed for assessing nuance-oriented reliability of LLMs in a systematic way.
Using IFEval++, we comprehensively evaluate 20 proprietary and 26 open-source LLMs. Our analysis reveals that current LLMs often exhibit considerable inconsistent competence across cousin prompts. The relative drop from accuracy on IFEval to reliable@10 on IFEval++ can be as large as 61.8% for Qwen3-0.6B and 54.7% for GPT-3.5-turbo-1106, while even the most reliable model, GPT-5, experiences a decrease of 18.3%. Notably, nuance-oriented reliability emerges as a second-order property: for instance, while Gemma-3-IT-27B ranks 17th in accuracy on IFEval, it rises to 7th when considering reliable@10 on IFEval++. Furthermore, we empirically characterize how models’ nuance-oriented reliability evolves with chronological development, model scale, reasoning capability, and augmentation type.
Observing the insufficiency of current LLMs in nuance-oriented reliability, we investigate three potential pathways for improvement: prediction-based methods, training-based methods, and test-time scaling. Among these, parallel test-time scaling through rejection sampling proves most effective, enabling a relatively weak model such as Qwen3-4B to surpass even the strongest open-source model, LLaMA-3.3-70B-Instruct.
We contend that improving this nuance-oriented reliability is a crucial next step toward dependable, trustworthy LLM behaviors. In summary, our contributions primarily lie in:
$\bullet$ We investigate the nuance-oriented reliability of LLMs in instruction-following and propose a new metric, reliable@k, to quantify it.
$\bullet$ We develop an automated pipeline for generating cousin prompts, and build IFEval++, a benchmark that enables actionable evaluation of nuance-oriented reliability.
$\bullet$ Using IFEval++, we conduct comprehensive experiments across a wide range of models and explore three distinct pathways to enhance reliability.
2 Preliminaries
2.1 Problem Statement
[…]
A Pilot Experiment. To ground our discussion, we begin with a pilot experiment focusing on test cases involving word-length constraints on responses. From IFEval, we extract 48 suitable prompts characterized by a relation (at least, at most, or around) and a numerical value num_word, as exemplified in Figure˜2. We then prompt GPT-5-mini to convert the prompt in each test case into a template, leaving the relation and value configurable. Using a length interval of 10, ranging from 50 to 800, we program to instantiate the templates, yielding 3×76×48=10,9443$\times$ 76$\times$ 48=10,944 prompts. The evaluation configurations are adjusted accordingly. As shown in Figure˜3, LLM performance is highly sensitive to the requested word count. A minor modification, e.g., from “at most 600” to “at most 610”, can incur many test cases to fail. This sensitivity is especially pronounced under the more challenging around constraint. This confirms that LLMs may not be reliable across prompts with subtle nuances. We ask: Shouldn’t a reliable LLM tackle them all?
Figure 3: Results of Requesting Varying Number of Response Words. Experiments with Qwen3-8B.
Two Reliability Dimensions in Instruction-Following. To situate our research, we review 36 benchmarks designed to evaluate instruction-following performance (see Section˜A.1). Most of these works focus on benchmark comprehensiveness—seeking evaluations that span diverse task types, domains, and constraint categories. This focus aligns with the view that a reliable assistant should handle diverse user demands across scenarios. However, our pilot study suggests that comprehensiveness alone is insufficient to capture the full notion of reliability. In real-world use, users may phrase identical requirements in multiple ways or express similar requirements with nuanced differences (see Figure˜2 for examples). A truly reliable LLM should consistently produce correct outputs across these prompts. We therefore propose a complementary perspective: nuance-oriented reliability, which measures the model’s stability when responding to “cousin prompts” that convey analogous user intents and differ only in subtle linguistic or semantic nuances.
2.2 Scope of This Work
[…]
Data Source & Constraint Types. We choose IFEval (Zhou et al., 2023) as our main testbed, since SOTA LLMs have already achieved strong performance on it (e.g., OpenAI’s GPT-5 reaches an accuracy of 95.9% as measured in Table˜1), making it a meaningful setting for discussing reliability and the challenging reliable@k metric. This benchmark focuses on single-turn tasks with 25 types of code-verifiable format constraints, such as “Answer using at most {num_words} words.” Reliably tackling such simple yet foundational instructions is a necessary condition for addressing more complex, multi-turn, and multi-modal instruction-following settings. See Section˜B.1 for the 25 constraints.
3 Curating Cousin Prompts At Scale
We introduce three data augmentation methods for producing cousin prompts (Section˜3.1), our checker for ensuring augmented test case validity (Section˜3.2), and the operational procedures for finalizing the IFEval++ benchmark (Section˜3.3).
Figure 4: Conceptual Illustration of the Augmentation-Filtering Data Curation Pipeline.
3.1 Data Augmentation Methods
Data augmentation for cousin prompts primarily involves revisions on two components: rewriting the prompts and adjusting the evaluation configurations, if necessary. In this work, we mainly account for three types of augmentation methods and employ LLM-based augmentation. See examples of the augmented cousin prompts in Figure˜2.
Type 1. Rephrasing. This reflects the scenario where different users may articulate the same task in varying ways. The augmenter rephrases each of the 541 prompts from IFEval into alternative wordings while preserving the core task and constraint semantics (see Section˜E.2). For rephrasing, the corresponding evaluation configurations remain unchanged, since the rephrased prompts still request responses that satisfy the same requirements as the original prompts.
Type 2. Distractor Addition. The augmenter appends each prompt with a distractor constraint that does not interfere with the satisfaction of the original constraints (see Section˜E.2). The augmenter is instructed to ensure that each distractor constraint remains compatible with the original prompt. In this setting, responses continue to be evaluated against the original constraints, with the added constraint functioning purely as a distractor, in contrast to prior work on multi-constraint composition (Jiang et al., 2024; Wen et al., 2024). This setup examines whether LLMs can integrate additional constraints in a plug-and-play manner.
Type 3. Constraint/Task Reconfiguration. In this setting, we aim to examine whether LLMs can reliably follow instructions when prompts are presented in varied scenarios. To this end, we craft cousin prompts that share the same requirement types but differ in subtle ways. For test cases with at least one configurable constraint (e.g., num_word), we instruct the augmenter to modify the constraint and update the evaluation configurations accordingly (see Section˜E.2). This mirrors the pilot experiment introduced in Section˜2.1. For cases with only non-configurable constraints, such as requiring responses in JSON format, we instead configure the task to indirectly impose the same requirement in a different context (Section˜E.2). In the case of task reconfiguration, the evaluation configurations remain unchanged.
3.2 Code-Assisted Validity Check
LLM-based augmentation enables us to rapidly acquire cousin prompts, but risks introducing flawed test cases. To mitigate this issue, we develop an automated validity checker.
The checker is designed with an emphasis on high recall: its primary objective is to flag any potentially erroneous cases, even at the cost of over-detection. This strategy allows us to curate a high-quality evaluation set by filtering candidates against the checker. Inspired by Chen et al. (2023), our central design is to embed the implementations of evaluation functions directly into the checker’s prompt, clearly specifying how the evaluation operates (see Section˜E.2). Empirically, this approach enables more deterministic and robust verification. The checker is implemented in a unified way to monitor all three types of augmented test cases.
Validating the Checker. Following principles in Section˜B.2, we iteratively refined the checker prompt on 3×300=9003$\times$ 300=900 test cases, where the augmenter was deliberately instructed to inject errors. This is an aggressive modeling of flawed prompts that the checker may encounter. Using GPT-5-mini (medium reasoning effort), the checker achieved a recall of 99.7% (898/900).
Table 1: Nuance-Oriented Reliability of Representative LLMs. We use rel@k for reliable@k and “C/T” for constraints or tasks. We report reliable@k across IFEval++ subsets by augmentation method. Models are ranked by reliable@10 on IFEval++ (with the rankings of IFEval accuracy for reference). Reasoning models are in blue, non-reasoning in beige. Best and second-best scores are in bold and underline.
Model IFEval +Rephrasing +Distractor +C/T Reconfig IFEval++
[…]
Open-Source LLMs
LLaMA-3.3-70B-Instruct 92.1 1 89.6 88.9 88.0 80.0 89.6 80.6 71.0 -22.9%
DeepSeek-V3.1 90.6 5 86.0 81.3 85.6 77.3 86.5 76.0 63.4 -30.0%
Qwen3-Next-80B-A3B-TK 91.3 2 86.1 80.8 84.8 76.5 86.3 76.9 63.2 -30.8%
Kimi-K2-0905 91.3 2 86.9 81.5 85.8 78.0 85.2 76.0 63.0 -31.0%
Qwen3-Next-80B-A3B-IT 90.9 4 84.6 80.4 86.5 77.3 86.5 75.8 62.3 -31.5%
GLM-4.5 86.9 10 82.6 78.2 81.1 72.5 83.2 74.9 61.7 -28.9%
Gemma-3-IT-27B 84.3 17 80.4 77.8 79.1 73.4 80.4 70.2 61.6 -27.0%
Qwen3-14B 88.4 6 85.4 81.5 83.2 73.4 83.2 72.5 61.2 -30.8%
Gemma-3-IT-12B 84.3 17 81.1 78.2 79.7 73.0 79.3 69.5 59.5 -29.4%
Qwen3-8B 87.6 7 83.4 79.5 81.7 72.8 81.3 69.9 58.8 -32.9%
LLaMA-3.1-70B-Instruct 85.6 15 82.3 79.1 79.5 71.0 81.0 69.1 57.1 -33.3%
Qwen3-30B-A3B 87.6 7 80.6 75.6 81.5 70.6 82.6 71.2 57.1 -34.8%
GLM-4.5-Air 87.4 9 82.6 76.3 81.1 70.8 81.5 70.8 56.7 -35.1%
Qwen3-235B-A22B 86.3 12 81.0 73.8 78.6 69.7 81.3 70.1 56.6 -34.5%
GPT-oss-120b 86.0 14 81.9 75.4 80.0 72.6 77.3 67.1 56.4 -34.4%
Qwen2.5-72B-Instruct 86.7 11 82.6 76.3 77.6 68.2 80.0 69.1 55.6 -35.9%
Qwen3-32B 86.3 12 81.5 75.8 81.9 72.3 81.7 69.7 55.3 -36.0%
DeepSeek-R1 83.4 20 77.1 70.6 77.8 69.1 75.8 65.4 53.0 -36.4%
Qwen3-4B 85.2 16 79.9 74.5 78.7 67.8 79.1 65.4 52.1 -38.8%
Qwen2.5-32B-Instruct 82.4 21 77.6 71.2 72.5 62.8 75.4 64.1 50.3 -39.0%
GPT-oss-20b 83.9 19 76.5 68.2 76.9 66.2 73.2 61.9 46.8 -44.3%
Qwen2.5-14B-Instruct 78.2 22 72.6 67.1 69.5 59.0 69.7 57.3 44.7 -42.8%
LLaMA-3.1-8B-Instruct 75.8 23 70.8 64.7 65.8 57.1 67.1 53.8 41.4 -45.4%
Qwen2.5-7B-Instruct 73.0 24 66.2 60.1 62.5 50.5 63.8 50.6 34.8 -52.3%
Qwen3-1.7B 71.5 25 63.4 55.1 62.3 49.4 61.6 46.2 34.0 -52.5%
Qwen3-0.6B 58.0 26 47.5 41.4 46.0 33.3 46.6 34.2 22.2 -61.8%
Proprietary LLMs
GPT-5 95.9 1 93.3 90.4 93.0 86.5 93.7 88.7 78.4 -18.3%
o3 94.3 3 93.5 89.1 92.1 84.5 91.7 85.0 75.0 -21.3%
o4-mini 95.6 2 91.9 87.1 90.4 81.7 91.9 84.8 72.3 -24.4%
Gemini-2.5-pro 93.3 5 89.6 87.4 88.4 82.6 91.3 83.0 71.9 -23.0%
GPT-5-mini 94.1 4 90.6 85.2 89.3 83.0 89.1 82.8 70.6 -25.0%
o3-mini 91.4 6 89.5 86.0 88.7 81.5 88.5 80.6 70.1 -25.2%
Gemini-2.5-flash 90.4 8 87.6 84.3 85.6 78.4 86.7 78.2 66.4 -26.5%
Gemini-2.0-flash 90.6 7 87.8 83.7 84.4 78.0 85.6 77.3 65.8 -27.3%
GPT-4.1 89.8 9 86.3 83.0 83.9 76.3 84.3 76.9 64.7 -28.0%
Gemini-2.5-flash-lite 88.0 10 83.9 81.5 83.2 75.8 83.0 73.2 62.7 -28.8%
GPT-5-nano 90.6 7 85.0 81.3 85.8 77.3 83.9 71.5 61.6 -32.0%
GPT-4o 85.0 13 80.2 76.3 79.7 71.3 79.3 69.1 58.6 -31.1%
GPT-4.1-mini 86.9 11 83.2 78.9 79.9 71.7 80.6 70.2 58.4 -32.8%
GPT-4o-2024-05-13 85.4 12 81.3 78.0 79.1 71.7 79.1 66.7 57.9 -32.3%
GPT-4o-2024-11-20 80.8 15 77.6 73.9 75.4 68.8 75.2 64.1 55.3 -31.6%
GPT-4o-mini 80.4 16 75.8 71.7 72.1 64.1 72.8 60.3 48.4 -39.8%
o1-mini 83.4 14 76.2 66.9 74.7 65.2 73.8 59.3 46.4 -44.3%
GPT-4.1-nano 77.4 17 72.5 67.8 70.1 60.6 68.8 55.6 42.9 -44.6%
GPT-3.5-turbo-0125 66.4 18 57.9 50.8 55.5 44.7 55.3 42.1 32.2 -51.5%
GPT-3.5-turbo-1106 61.6 19 53.0 48.4 50.5 41.0 51.8 37.8 27.9 -54.7%
To further assess robustness, we evaluated it on 3×1,000=3,0003$\times$ 1,000=3,000 additional flawed cases, maintaining a recall of 99.9% (2,997/3,000); the few misses were due to unavoidable hallucinations (Sun et al., 2025). The checker also achieved 90.0% precision on 541 unrevised promptsMany flagged cases were indeed flawed or ambiguous. Necessary corrections are detailed in Section B.3, and the revised dataset is used for subsequent experiments. . Overall, the checker rigorously detects invalid prompts with minimal valid loss. Moreover, since the augmenter is also explicitly guided to avoid introducing flaws, the two-layer mechanism minimizes the risk of flaws arising from the augmentation process.
3.3 Synthesizing the IFEval++ Benchmark
In this section, we synthesize IFEval++: To balance cost and coverage, each original test case is expanded into nine replicates, with three from each augmentation method. Following the augmentation methodology introduced earlier, we automatically produce cousin prompts and verify them using the code-assisted validity checker to ensure integrity. This process is iterated until a sufficient number of valid cousin prompts are collected.
The IFEval++ benchmark comprises 541 test cases, each containing one original prompt and nine cousin prompts. With IFEval++, we can characterize the nuance-oriented dimension of LLMs’ reliability, which cannot be revealed by evaluating only standalone prompts. Accordingly, we report reliable@1 (i.e., accuracy on IFEval), reliable@10 on IFEval++, and reliable@4 and reliable@2 on the subsets of IFEval++ corresponding to each augmentation type.
4 Experiments
4.1 Experimental Setup
Models. In this work, we include a diverse set of 46 models to facilitate comparative analyses across four key dimensions: (a) differences in model scale, (b) chronologically different generations of models, (c) contrasts between open-source and proprietary models, and (d) distinctions between reasoning and non-reasoning models. Additional model details are provided in Appendix˜C.
Generation Strategies. We use official default system messages when provided in chat template (e.g., for Qwen-2.5 models); otherwise, we leave the system message empty to elicit the model’s default behavior, as suggested by models like DeepSeek-R1https://huggingface.co/deepseek-ai/DeepSeek-R1#usage-recommendations. We further validate this choice in Section˜D.1. We employ greedy decoding by default, which is suitable for reliability evaluation and provides a reproducibility guarantee. Additionally, we enable reasoning mode by default if supported.
Evaluation Methods. Following the practices of DeepSeek-V3 (Liu et al., 2024), Qwen3 (Yang et al., 2025a), and Kimi K2 (Team et al., 2025), we use the prompt strict evaluation mode of IFEval. This requires LLMs to satisfy all requirements specified in the prompts. Based on this success signal, we compute the reliable@k metric.
4.2 Main Results with IFEval++
We answer the initial question: Current LLMs do not guarantee nuance-oriented reliability.
As shown in Table˜1, the relative drop from accuracy on IFEval to reliable@10 on IFEval++ is substantial, reaching up to 61.8% for Qwen3-0.6B and 54.7% for GPT-3.5-turbo-1106. Even the most reliable model, GPT-5, suffers an 18.3% decline when confronted with prompts containing subtle variations. Importantly, we observe that higher accuracy on IFEval does not necessarily translate into higher reliable@10 on IFEval++. For example, while Gemma-3-IT-27B ranks only 17th on IFEval, it rises to 7th place on IFEval++. This highlights that nuance-oriented reliability is a second-order property beyond accuracy, one that comprehensiveness-oriented benchmarks fail to capture. We provide the category-wise reliability in Section˜D.4. Besides, we arrive at several empirical findings:
$\bullet$ Chronological Development: As expected, more recent LLMs tend to perform better, especially when developed by the same vendor. This underscores the importance of the underlying training methodology, as highlighted by the stark difference in reliability between LLaMA-3.3-70B-Instruct (reliable@10: 71.0) and its predecessor, LLaMA-3.1-70B-Instruct (reliable@10: 57.1).
$\bullet$ Model Scale: In general, larger-scale models tend to outperform smaller ones. An exception is observed within the Qwen3 family, where Qwen3-14B demonstrates lower accuracy on IFEval but achieves a higher reliable@10 score compared to the larger Qwen3-32B. This indicates that surpassing a certain parameter size is sufficient to attain high reliability. Meanwhile, performance is not determined by scale alone: Training data quality and methodology also play crucial roles. Notably, models with comparable parameter counts but developed by different vendors can show substantial differences in reliability, as exemplified by Gemma-3-IT-27B and Qwen2.5-32B-Instruct.
$\bullet$ Reasoning: Reasoning models demonstrate increasingly stronger nuance-oriented reliability. However, LLaMA-3.3-70B-Instruct, though not a reasoning model, still achieves the highest ranking among open-source models. This suggests reasoning itself may not be a prerequisite for achieving high reliability. This motivates us to examine the specific contribution of reasoning to reliability, which we investigate in Section˜5.3.
$\bullet$ Augmentation Type. Models exhibit varying robustness to the three types of cousin prompts. Most models can effectively handle rephrased cousin prompts, which share the greatest similarity with the original ones. This result is expected, as rephrased prompts differ only in wording (understanding the instructions) while requesting the same underlying response (executing the instructions). In contrast, cousin prompts that include additional distractor constraints or C/T reconfiguration introduce greater complexity. Their corresponding reliable@4 scores are typically lower than those of rephrased prompts. We hypothesize that, although distractor constraints are designed to remain compatible with the original instructions, and C/T reconfiguration introduces only slight variations in the underlying requirements, both inevitably increase the difficulty of response planning and execution (Dong et al., 2025b).
Table 2: AUROC Scores of Prediction Methods.
Method Qwen3-8B Qwen2.5-7B
Verb. Conf. 0.549 0.518
Perplexity 0.497 0.529
Probing 0.757 0.759
5 How to Improve Reliability?
While LLMs are not yet reliable, we explore three possible directions for improving them.
5.1 Predicting the Selective Following
If we can predict how the LLM reacts to different prompts, we can proactively approach reliable LLM services by prompt selection. This case study goes in a similar vein to Stolfo et al. (2025); Cao et al. (2024); Heo et al. (2025a, b), but our focus is to predict the instruction-following prior to generation and across various cousin prompts. We experiment with Qwen3-8B and Qwen2.5-7B-Instruct. For each model, we randomly sample 1,200 of them as the validation set, with 600 followed and 600 unfollowed. We evaluate three prediction methods, with performance presented in Table˜2.
$\bullet$ Verbalized Confidence. We instruct the LLM to articulate its confidence in successfully following one given prompt, on a scale from 0 to 9. Yet, the self-reported confidence yields performance close to random guessing. This outcome can be attributed to the LLMs’ general tendency toward overconfidence, aligning with Xiong et al. (2024).
$\bullet$ Prompt Perplexity. We compute the perplexity of LLMs on each prompt and use it for predicting instruction-following. Yet, the close-to-chance AUROC reveals that low-perplexity prompts that are most familiar to the model do not necessarily correspond to successful instruction-following.
$\bullet$ Probing Hidden States. Causal LMs may plan their responses before generation, and the relevant signals can be probed from their hidden states (Dong et al., 2025b). Specifically, we train a linear logistic regression probe, which takes the hidden states of the last input token as input. To train the probe, we additionally collect a non-overlapping set of 200 followed and 200 unfollowed prompts. Compared with the previous two methods, probing provides a more effective prediction. The best validation performance is achieved when probing the 17th and the 27th layer on two models, respectively. Although probing shows potential in this direction, the probes are far from being a reliable predictor.
Figure 5: Training on Datasets Affects Reliability Performance on IFEval++ in Varying Fashion.
5.2 Training-Based Method
We adopt supervised fine-tuning (SFT) as our training method and evaluate two dataset configurations, which represent the two extremes of prompt relatedness. (1) We use the general-domain instruction-following dataset, alpaca (Taori et al., 2023), whose prompts are largely unrelated to IFEval++. (2) We curate an additional set of prompts, also augmented from IFEval but decontaminated from IFEval++ prompts via exact matching. Then, we collect responses via rejection sampling against LLaMA-3.3-70B-Instruct. We fine-tune Qwen-2.5-7B-Instruct for 312 steps on each dataset. Further details on data curation and training are provided in Section˜E.1.
Figure˜5 reports the reliability performance on IFEval++ across training steps. The two datasets lead to markedly different outcomes. Models fine-tuned on alpaca exhibit a slightly declining performance. In contrast, training on the curated cousin prompts steadily improves reliability, with performance climbing above 45% after 200 steps. This reaffirms reliability as a second-order property from the training perspective. It benefits more from targeted fine-tuning on semantically adjacent samples rather than from the sheer scale of training data. This observation resonates with recent explorations on rewriting training data (Fujii et al., 2025; Team et al., 2025). In Section˜E.2, we additionally explore how different training techniques (SFT vs. DPO (Rafailov et al., 2024)) affect the nuance-oriented reliability.
Figure 6: Impact of Reasoning Effort on Nuance-Oriented Reliability of LLMs.
5.3 Test-Time Scaling
We explore two forms of scaling up test-time compute (Snell et al., 2024): sequential and parallel.
$\bullet$ Sequential (Reasoning Effort). Reasoning models inherently capture thinking patterns such as self-correction (Guo et al., 2025). Scaling up reasoning effort can be viewed as a sequential analogue to scaling up test-time compute. We include four models that support adaptive reasoning budget control: GPT-5-mini, Gemini-2.5-Flash, Qwen3-14B, and Qwen3-32B. As shown in Figure˜6, increasing reasoning effort usually improves reliability, particularly for GPT-5-mini. Manual inspection of their intrinsic CoT reveals that, through reasoning, models can infer user intent and adapt responses accordingly. This explicit reasoning process enhances robustness against subtle prompt variations. However, inappropriate reasoning patterns may reduce reliability (Chen et al., 2024a; Dong et al., 2025a), as evidenced by the negative effects observed in Qwen3-32B.
$\bullet$ Parallel (Rejection Sampling). To scale test-time compute in parallel, we adopt sampling rather than greedy decoding, generating nn samples with a temperature of 1. We assume the availability of a response selector. For example, in prompts related to IFEval++, format requirements can be extracted with the assistance of an LLM, while Python code can be used for rapid response evaluation. With the response selector, the reliable@10 metric is scored positively if the 10 cousin prompts within one group all have at least one response that passes among the nn samples. The results are shown in Figure˜7. Rejection sampling with a suitable response selector effectively enhances the reliability of LLMs in instruction-following. The reliable@10 scores rise significantly as more samples are allowed, plateauing around n=12n=12. We further observe that reasoning models, such as Qwen3-8B, benefit more from additional compute via parallel scaling, given their more exploratory nature (Chen et al., 2025). For instance, Qwen3-4B and Qwen3-8B, with only n=3n=3 samples, already surpass LLaMA-3.3-70B (the strongest open-source model in Table˜1). However, we note that extending this parallel scaling strategy to other domains may be challenging without suitable response selectors.
Figure 7: Scaling Landscape of Rejection Sampling.
6 Conclusion
In this work, we present a systematic study on the nuance-oriented reliability of LLMs in instruction-following. To characterize this property, we introduce the new metric reliable@k, which can be instantiated with cousin prompts. We design and implement a fully automated pipeline to generate such cousin prompts through data augmentation methods, combined with a code-assisted validity checker. Leveraging this pipeline, we construct the IFEval++ benchmark, which enables systematic evaluation. Extensive experiments across 46 models using IFEval++ uncover a critical vulnerability: current LLMs frequently fail to ensure reliability across cousin prompts with nuanced differences. We further explore three directions for improving this reliability: prediction-based, training-based, and test-time scaling. Overall, our findings highlight the importance of moving toward nuance-oriented reliability as a key dimension of dependable, trustworthy LLM behaviors.
7 Acknowledgment
This work was supported by Ant Group and the Center for High Performance Computing at Tsinghua University.
8 Limitations
We identify several points that are not fully solved in this work and are worth future investigation.
Measurement Efficiency. Our practice of evaluating nuance-oriented reliability is straightforward, where we introduce cousin prompts to avoid the risk of hacking certain test cases. This enables us to compute the reliable@k metric to monitor the reliability dimension. However, this also introduces a higher evaluation cost. For example, a full evaluation of IFEval++ requires 10× as many responses to be generated as IFEval. Future work will focus on selecting or crafting those most differentiating test cases and cousin prompts, which enables a more cost-efficient evaluation.
Missing Content-Level Evaluation. Following common practice in prior studies, we primarily evaluate whether LLMs adhere to constraint instructions. Moreover, prompts often include one or multiple tasks, which are not explicitly assessed by either IFEval or IFEval++. This design choice is acceptable, as satisfying constraints is a necessary prerequisite for correctly following the entire prompt. Future work could extend this by jointly measuring content-level response quality along with constraint adherence.
Focus on IFEval. In this work, we primarily adopt IFEval as our testbed. It is widely recognized in the community as a standard benchmark for instruction-following evaluation and has approached performance saturation with the latest models. Nevertheless, our reliability evaluation methodology is not limited to this single benchmark. Future research may extend our approach to other existing benchmarks, or even beyond the instruction-following domain. We envision that evaluating nuance-oriented reliability will become a default practice, complementing traditional capacity-oriented evaluation.
Potential Bias in Automated Validity Check. Although the code-assisted validity checker demonstrates high recall and precision, its automation depends on LLM-based judgment, which may introduce subtle biases or misclassifications. We carefully and systematically validate the checker’s performance in Section˜3.2 and additionally dedicate substantial, though undocumented, manual effort to continuously monitor and ensure the quality of the checking process. Moreover, the strong performance of advanced models—such as GPT-5, which achieves a reliable@10 of 78.4—suggests that most test cases are indeed valid and feasible; otherwise, invalid cases would cause even the strongest LLMs to fail consistently.
Limited Coverage of Improvement Recipes. While numerous efforts have been proposed to enhance instruction-following capabilities, reproducing all such methods to assess their combined impact on nuance-oriented reliability lies beyond the scope of this work. We focus on several representative improvement strategies, including prediction-based, training-based, and test-time scaling approaches. Future research could undertake a more comprehensive investigation to identify the most effective techniques for improving reliability across diverse enhancement paradigms.
Cross-Lingual Reliability. Our study primarily evaluates instructions expressed in English, leaving the applicability of reliable@k to under-resourced languages largely unexplored. In principle, the metric can be adapted through practical adjustments, such as incorporating a translation module in the data augmentation pipeline and extending evaluation functions to additional languages. Assessing reliability across languages expands the scope of model evaluation and may reveal critical insights into the fairness and generalizability of LLM performance for users of diverse linguistic backgrounds. We envision that a comprehensive investigation of cross-lingual reliability will yield further insights into model generalization.
9 Ethical Considerations
Our research adheres to the ethical guidelines established by the Association for Computational Linguistics (ACL)§§§https://aclrollingreview.org/responsibleNLPresearch/. We have made every effort to conduct this work with the utmost respect for ethical principles and research integrity.
This research aims to advance the understanding of the reliability of large language models in instruction-following. All experiments were performed on publicly available models and datasets, ensuring compliance with relevant terms of service and data usage policies, e.g., IFEval (Zhou et al., 2023). No proprietary or confidential information was accessed or reverse-engineered during this study. Our analyses do not involve human subjects, sensitive personal data, or the generation of harmful content. Our use of AI assistants is limited to writing polishing and grammar checking. We release the complete codebase and datasets (e.g., IFEval++) to ensure full reproducibility of our results.
Finally, we remind the readers that any techniques introduced in this paper should be applied ethically and within appropriate research contexts.
Appendix A Related Works
A.1 Benchmarking Instruction-Following
Table 3: Key Designs of the Existing Instruction-Following Benchmarks.
Benchmark # Cases Comprehensiveness-Oriented Nuance-Oriented
New Tasks New Constraints Complicated Scen.
Collie-v1 (Yao et al., 2024) 2,080
CELLO (He et al., 2024a) 523
NPB (Sun et al., 2023)
FollowBench (Jiang et al., 2024) 820
IFEval (Zhou et al., 2023) 541
CoDI-Eval (Chen et al., 2024c) 9,060
InFoBench (Qin et al., 2024) 500
FOFO (Xia et al., 2024) 494
AlpacaEval-LI (Yuan et al., 2024) 802
MT-Bench-LI (Yuan et al., 2024) 240
SIFo (Chen et al., 2024b) 800
RuleBench (Sun et al., 2024)
ComplexBench (Wen et al., 2024) 1,150
CFBench (Zhang et al., 2024b) 1,000
SysBench (Qin et al., 2025) 500
MMMT-IF (Epstein et al., 2024) 71
RealInstruct (Ferraz et al., 2024) 302
Multi-IF (He et al., 2024b) 4,501
LIFBench (Wu et al., 2025) 2,766
VFF (Wang et al., 2025b) 21k
StructFlowBench (Li et al., 2025a) 155
PBIF (Zeng et al., 2025) 24k
CodeIF (Yan et al., 2025) 1,200
WildIFEval (Lior et al., 2025) 11,813
XIFBench (Li et al., 2025b) 465
CodeIF-Bench (Wang et al., 2025a)
MulDimIF Ye et al. (2025) 1,200
MathIF (Fu et al., 2025) 420
AGENTIF (Qi et al., 2025) 707
LIFEBench (Zhang et al., 2025b) 10,800
MaXIFE (Liu et al., 2025) 18,285
EifBench (Zou et al., 2025) 1,000
Ordered CommonGen (Song et al., 2025) 27,648
IFBench (Pyatkin et al., 2025) 300
IFScale (Jaroslawicz et al., 2025)
IFEvalCode (Yang et al., 2025b) 1,600
We provide a chronological survey in Table˜3.
Collie-v1 (Yao et al., 2024). This work proposes a grammar-based framework for constructing controlled text generation, allowing automatic conversion from formal constraint configurations to natural language instructions. The final compiled benchmark includes 2,080 instances comprising 13 constraint structures.
CELLO (He et al., 2024a). Accounting for the complexity of in-the-wild user instructions, this work focuses on long and complex instructions. The proposed CELLO benchmark contains 523 instructions that cover four types of response constraints: count limit, answer format, input dependency, and phrase callback.
NPB (Sun et al., 2023). This work evaluates the controllability of LLMs during generation tasks. It includes numerical planning constraints, requesting responses with varying lengths.
FollowBench (Jiang et al., 2024). This work assesses the constraint-following abilities of LLMs, considering five types of constraints: content, situation, style, format, and example. It involves incrementally adding constraints when curating instructions, yielding a total of 820 instructions with different difficulty levels.
IFEval (Zhou et al., 2023). This work pioneers in proposing a broad taxonomy of code-verifiable constraints. The resulting benchmark consists of 541 instructions, covering 25 types of constraints.
CoDI-Eval (Chen et al., 2024c). This work introduces expansion and rewriting steps to enhance instruction diversity. The resulting CoDI-Eval benchmark consists of 9,060 instructions and covers a range of constraints, including sentiment, topic, keyword, length, and toxicity avoidance.
InFoBench (Qin et al., 2024). This work proposes a novel scoring method: preparing a checklist for each instruction and employing LLM-as-a-Judge to decide the acceptance. Meanwhile, it introduces the InFoBench benchmark, which comprises 500 instructions and involves constraints like content constraint, linguistic constraint, style rules, format specifications, and number limits.
FOFO (Xia et al., 2024). This work specifically studies the format-following abilities of LLMs. It proposes the FoFo benchmark, which contains 494 instructions. Each test case requires the responses to adhere to domain-specific formats, e.g., medical reports and Latex. LLM-as-a-Judge is employed for automated adherence evaluation.
AlpacaEval-LI & MT-Bench-LI (Yuan et al., 2024). This work investigates LLMs’ ability to follow length constraints in instructions. To facilitate this evaluation, the authors extend AlpacaEval (Dubois et al., 2024) and MT-Bench (Zheng et al., 2023) by constructing the corresponding length-instructed variants
SIFo (Chen et al., 2024b). This work focuses on a specific aspect of instruction-following evaluation—ordering of multiple instructions (i.e., constraints in this work). The SIFo benchmark comprises 800 test cases, with LLMs explicitly instructed to address multiple constraints in a prescribed sequential order. This is implemented as a single-turn interaction.
RuleBench (Sun et al., 2024). This work studies inferential rule following, where additional if–then rules are incorporated into the instructions. This setting differs from the affirmatively requested constraints in other works.
ComplexBench (Wen et al., 2024). This work studies LLMs’ abilities to follow complex instructions involving multiple constraint compositions. The ComplexBench consists of 1,150 instructions, each associated with a constraint composite. The composition types considered include And, Chain, Selection, and Nested.
CFBench (Zhang et al., 2024b). This work emphasizes the comprehensiveness of constraint types for evaluating instruction following. It proposes a constraint system with 25 subcategories and meticulously curates 1K Chinese instructions.
SysBench (Qin et al., 2025). This work echoes the real-world user practices involving system prompts. They evaluate the instruction-following ability of LLMs, involving 500 system prompts requesting six types of constraints.
MMMT-IF (Epstein et al., 2024). This work extends the instruction-following evaluation to multi-modal and multi-turn settings. The MMMT-IF benchmark contains 71 test cases, with image inputs and multiple constraints scattered across chat.
RealInstruct (Ferraz et al., 2024). This paper introduces real-world user queries as a source for instruction-following test cases. The authors meticulously curate instruction candidates from ShareGPT and then decompose them into tasks, contexts, and constraints. The final test set includes 302 instances, each of which typically contains multiple constraints.
Multi-IF (He et al., 2024b). This work extends instruction-following evaluation to multi-turn and multilingual settings. The Multi-IF benchmark is composed of 4,501 multilingual conversations, constructed by augmenting IFEval with multi-turn interactions and translations.
LIFBench (Wu et al., 2025). This work studies instruction-following abilities of LLMs in long-context tasks, e.g., processing multiple documents. The benchmark consists of 2,766 instructions, involving long-context tasks and 6 types of instruction-following capabilities.
VFF (Wang et al., 2025b). This work exclusively explores the format-constraint following abilities of LLMs. It constructs 21K test cases with varying levels of difficulty by combining one to three verifiable constraints per instruction. The constraint types include limited word count, limited content, limited punctuation, limited structure, limited grammar, and specific number formatting.
StructFlowBench (Li et al., 2025a). This work explores the instruction-following performance of LLMs in multi-turn scenarios, emphasizing the structural dependencies that arise across turns. The benchmark incorporates eight types of constraints and six categories of structural flows: follow-up, refinement, recall, expansion, summary, and unrelatedness. In total, it includes 155 dialogues, evaluated via LLM-as-a-Judge.
PBIF Zeng et al. (2025). This work investigates how the ordering of multiple constraints influences the instruction-following performance of LLMs. It identifies an intriguing phenomenon: LLMs often perform better when the constraints are presented in a hard-to-easy order.
CodeIF (Yan et al., 2025). This work introduces a benchmark for evaluating LLMs’ instruction-following in code generation, covering tasks like synthesis, debugging, refactoring, and explanation. It comprises 1,200 test cases across 8 categories and 50 sub-instructions, spanning multiple programming languages.
WildIFEval (Lior et al., 2025). This work curates a set of 12K complex instructions collected from real-world user queries on Chatbot Arena (Chiang et al., 2024). Each instruction is then decomposed into 8 categories of constraints, and LLM-as-a-Judge is employed for automated evaluation of instruction-following.
XIFBench (Li et al., 2025b). This work primarily explores the instruction-following abilities of LLMs across different language settings, from high-resource (English and Chinese) to low-resource (Hindi). It starts with five categories of constraints: content, style, situation, format, and numerical. By composing between one and five constraints into each seed instruction, the authors construct a total of 465 hard instructions. These are further extended into multiple languages through translation, resulting in a multilingual benchmark.
CodeIF-Bench (Wang et al., 2025a). This work focuses on multi-turn interactive code generation. It covers 9 types of verifiable instructions aligned with real-world software development requirements for test-case validation, and supports both Static and Dynamic Conversation settings.
MulDimIF Ye et al. (2025). This work proposes a multi-dimensional constraint framework, involving three types of constraint formatting patterns (Example, Listing, and Incorporation), four categories of constraints, and four difficulty levels. In total, this benchmark consists of 1,200 code-verifiable instructions.
MathIF (Fu et al., 2025) is a dedicated benchmark for evaluating instruction-following in mathematical reasoning tasks. It contains 420 evaluation samples derived from 15 Python-verifiable constraints across 4 categories, applied to math problems of varying difficulty.
AGENTIF (Qi et al., 2025) evaluates LLMs’ instruction-following abilities in agentic scenarios with long, complex, and constraint-heavy instructions. It comprises 707 human-annotated instructions drawn from 50 real-world agentic tasks.
LIFEBench (Zhang et al., 2025b). This work is conducted with a focus on length constraints, ranging from 16 to 8192 words. It contains 10,800 instructions, in English or Chinese, requesting responses with length satisfying three types of relations (At Most, At Least, and Equal To).
MaXIFE (Liu et al., 2025). This work focuses on multilingual and cross-lingual scenarios. It consists of parallel data in 23 languages, with each record combining a Basic Question and 1–3 Instructions, totaling 18,285 entries.
[…]
Ordered CommonGen (Song et al., 2025) introduces a benchmark that evaluates LLMs’ ability to generate concepts in a specified order. It contains 27,648 instances from 192 concept sets and 6 instruction templates, using ordered coverage to measure adherence to the specified sequence.
IFBench (Pyatkin et al., 2025). This work highlights the limited generalization of LLMs’ instruction-following abilities across constraint types, thus falling into comprehensiveness-oriented reliability. They curate a new set of 300 instructions, involving 58 new constraint types beyond the 25 types in IFEval.
IFScale (Jaroslawicz et al., 2025). This work introduces a benchmark for evaluating instruction-following performance as instruction density increases in professional business report generation. Each task contains 10–500 keyword-inclusion instructions drawn from a curated 500-term business vocabulary, with density increasing in steps of 10.
IFEvalCode (Yang et al., 2025b) introduces a multilingual benchmark for controlled code generation, evaluating LLMs’ ability to follow instructions beyond correctness in coding scenarios. It includes 1,600 samples in 8 languages with paired Chinese–English queries, assessing both correctness and instruction adherence.
Among these, several works are relevant to the techniques and the research problem studied in this work. We distinguish ours from several highly relevant ones:
  • $\bullet$
    CoDI-Eval (Chen et al., 2024c): This benchmark mainly focuses on the controllability of LLMs in generation tasks. This work advocates an instruction diversification process to synthesize diverse forms of constraint expression (i.e., the rewriting step). They aim to achieve finer-grained coverage of instructions rather than to study the reliability issue. What’s more, they fail to account for inter-prompt relationships.
  • $\bullet$
    SIFo (Chen et al., 2024b) & PBIF (Zeng et al., 2025): These two works investigate the position bias problem of LLMs in the context of instruction-following. They examine how LLMs’ compliance changes when the order of constraints within prompts is altered. This phenomenon can be viewed as a special case of our rephrasing-based augmentation approach.
  • $\bullet$
    LIFEBench (Zhang et al., 2025b): This work specifically studies length constraints. Their study is analogous to our pilot study in Section˜2.1. However, they focus on how long-context LLMs follow length constraints rather than on the general reliability.
  • $\bullet$
    IFScale (Jaroslawicz et al., 2025): This work studies how many instructions LLMs can simultaneously follow. For the research question, the study also relates to reliability, but it stress-tests LLMs along the complexity-oriented dimension.
A.2 Evaluating LLMs’ Reliability
There exist works revealing LLMs’ sensitivity to prompt wordings (Sclar et al., 2024). Cao et al. (2024) specifically study the worst performance across semantically equivalent case-level queries. Mizrahi et al. (2024) advocate for multi-prompt LLM evaluation, which shares the same methodology as our curation of cousin prompts. Based on these works, we move beyond studying simply prompt sensitivity issues (the rephrasing augmentation may reflect) and towards more general nuance-oriented reliability.
A critical distinction must be drawn between the reliable@k metric and sampling responses for one prompt independently kk times. Yao et al. (2025) introduce a new metric, passˆk, for evaluating agentic scenarios. While passˆk measures reliability across kk independent trials using the same prompt (often with temperature >0>0), our reliable@k metric evaluates consistency across distinct cousin prompts with nuanced variations. To quantitatively compare these two dimensions, we conducted an experiment sampling each IFEval prompt 10 times at temperature 1 to compute passˆ10, contrasting it with our reliable@k on IFEval++.
Table 4: Comparison between standard Accuracy, reliable@10, and passˆ10.
Model Name Accuracy reliable@10 passˆ10
(IFEval) (IFEval++) (Repeated Sampling)
Qwen2.5-7B-Instruct 73.0 34.8 53.0
Qwen3-4B 85.2 52.5 67.0
Qwen3-8B 87.6 58.8 71.0
Llama-3.3-70B-Instruct 92.1 71.0 85.6
As shown in Table˜4, reliable@10 and passˆ10 degrade at different rates compared to standard accuracy. For instance, even for strong models like LLaMA-3.3-70B, the gap between single-prompt accuracy (92.1) and cousin-prompt reliability (71.0) is significant, and distinct from the gap caused by sampling variance (85.6). This highlights that they capture different aspects of LLMs in instruction-following: passˆk reflects stability against random seed noise, whereas reliable@k reflects robustness against semantic nuances. Therefore, the two evaluation methods are orthogonal and can be jointly used to characterize the lower bound of an LLM’s reliability more rigorously. Furthermore, as we discuss in Section˜5.3, the variability exposed by multiple samples (whether from repeated sampling or cousin prompts) can be exploited via rejection sampling to improve final performance.
Appendix B Data Adaption
B.1 Constraint Types
We adopt the 541 test cases from IFEval. All of them are composed of tasks with several format constraints. The format constraints are categorized into 25 types, which we detail in Table˜6. These simple constraints lay the foundation for LLMs’ solving various more complex requirements. As most LLMs, as reported by the community and verified by our experiments (the ACC column in Table˜1), have boosted performance in the original IFEval, it is a necessary next step to understand how solid and reliable their excellent instruction-following abilities are. This motivates the nuance-oriented reliability examination.
B.2 Principles for High-Quality Data
Recall that each test case mainly relates to three key components: the prompt which instructs LLMs, the evaluation function for checking the instruction-following behaviors (specifically, Python code and response formats in the context of IFEval and IFEval++), and the evaluation configurations which specify how the evaluation functions should examine the instruction-following behaviors.
We document four principles for high-quality test cases, which are also enforced in our data augmentation in Section˜3.1 and implemented in our automated validity checker (see Section˜E.2).
  • $\bullet$
    Clear Instruction Intents: Each test case must contain an unambiguous and self-contained prompt that clearly specifies the task to be performed and the constraints to be satisfied. The prompt should minimize interpretation uncertainty and avoid overlapping objectives (e.g., mixing summarization with classification). This ensures that the model’s output can be attributed to understanding and following the instruction rather than guessing the task’s intent.
  • $\bullet$
    Verifiability through Evaluation Implementations: A high-quality test case must be verifiable by automated evaluation functions. The expected outputs or behavioral criteria should be formulated so that they can be programmatically validated without requiring human judgment. Since certain evaluation function implementations in IFEval possess inherent assumptions and inductive biases, the specified constraint requirements must remain within the boundaries of these evaluation functions. Meanwhile, the prompts must articulate the constraint requirements in a way that the downstream evaluation functions can verify them.
  • $\bullet$
    Alignment with Evaluation Configurations: The evaluation configuration (e.g., the relation and num_words for the length constraints) must be compatible with both the prompt and the evaluation function. Misalignments between prompt requirements and evaluation logic can produce false positives or negatives, undermining the reliability of the benchmark. Thus, each test case should be explicitly validated against the intended constraint requirements (i.e., the evaluation configurations) to ensure consistency.
  • $\bullet$
    Attainability: A test case should be solvable by a capable model based solely on the given prompt. Consider a scenario where two length constraints coexist within a single test case: one requiring more than ten sentences and another requiring fewer than eight words. Such constraints are inherently incompatible, rendering the test case infeasible and therefore subject to removal. We observe that IFEval includes several test cases exhibiting these attainability issues. Moreover, since our data augmentation strategies—namely, adding distractors and altering constraints—may occasionally combine multiple constraints, we design and implement our validity checker to carefully filter out any combinations that result in unattainable or contradictory requirements.
B.3 Revising Infeasible Test Cases
Manual review and community feedback reveal that a small number of test cases in IFEval are flawed. We make necessary yet minimal revisions to ensure test case integrity. The overall information and examples are listed in Table˜7. We identify that common issues in IFEval can be categorized into the following four classes.
Flawed Implementation of Evaluation Functions. A representative issue arises from mismatches between prompts and the underlying automatic checkers. For example, one test case requires the “#” character to appear at least four times. However, the original evaluation function only supports letters a-z, making it impossible to evaluate such symbols. As a result, this prompt is invalid for the evaluation function, despite being well-formed. We updated the evaluation function to accept any character—including symbols and digits—so that constraints targeting non-alphabetic characters can be correctly evaluated, ensuring the evaluation function aligns with the intended instructions.
Flawed Documentation of Evaluation Configurations. In some instances, the natural language prompt directly contradicts the evaluation arguments. For example, one prompt requires “the letter o to appear at least 6 times”, but the evaluation argument specifies “less than 6 times. Such inconsistencies lead to an unavoidable conflict between what the prompt demanded and how correctness is assessed. We corrected these cases by rewriting the prompt to match the evaluation arguments.
Typos in Natural Language Prompts. Some test cases contain minor typographical errors that make them misleading. For instance, one prompt instructs to separate two tables using “exactly 6 asterisk symbols: *******”, where the example actually has 7 asterisks, contradicting the requirement. This was revised to six asterisks “******” to match the stated constraint, thereby eliminating ambiguity introduced by typographical mistakes.
Unachievable Multi-Constraint Test Cases. Some prompts impose multiple constraints that are logically incompatible. A notable example requires the model first to repeat a sentence verbatim (including commas) and then simultaneously adhere to the instruction “avoid using commas”. Since the two conditions could not be satisfied simultaneously, such prompts were fundamentally unachievable. We resolved this by removing the contradictory requirement, keeping the prompt feasible.
Appendix C Model Details
In this work, we include a total of 36 large language models to establish a representative benchmark. Our selection aims to comprehensively capture the development trends of LLMs in terms of reliability. To this end, we include models that are widely recognized for their performance or influence within the research and developer communities, covering a diverse range of providers, architectures, and scales. For those proprietary models, we access their LLM services via API, specifically the OpenAI-compatible interface. For the open-source ones, we locally deploy them using vLLM (Kwon et al., 2023) for inference. We use the instruction-tuned or chat version and adopt official chat templates. See Table˜8 for the detailed information.
Figure 8: Results of Requesting Varying Number of Response Words. Experiments with Qwen3-32B.
Appendix D More Experimental Results
D.1 Impact of the System Prompt
Certain LLMs are published with default system prompts. For instance, models in the Qwen2.5 series include default prompts in their chat templates. In contrast, more recent reasoning models, such as DeepSeek-R1 and Qwen3, do not specify system prompts; their providers even recommend using an empty prompt. To elicit the default (and arguably optimal) behaviors while avoiding evaluation bias, our main experiments adopt the system prompt settings recommended by model providers.
Meanwhile, we notice that it is a common practice for service providers or end users to employ system prompts tailored for personalized LLM services (Ge et al., 2024; Qin et al., 2025). We are also curious about the impact of system prompt setting on the models’ reliability in instruction-following. We include six settings as listed in Table˜9: the default, three persons sampled from Ge et al. (2024), python coder, and encouraging better instruction-following. Except for the system prompt setting, other decoding strategies remain consistent.
The results, illustrated in Figure˜10, show that system prompts have an implicit but notable impact on performance in IFEval++. Empirically, the default system prompt tends to yield the best instruction-following performance, confirming the representativeness of our choice in the main experiments. In contrast, personalization may harm reliability: certain persona prompts reduce reliable@10 by up to 8.2%. Moreover, explicitly encouraging better instruction-following (the follow-instructions prompt) does not improve performance. Importantly, the same prompt may affect models differently—the persona3 prompt significantly degrades the reliability of Qwen3-8B, but has little or no impact on LLaMA-3.3-70B.
This initial empirical measurement substantiates the validity of our choice of using the default system prompt. However, the mechanisms by which the setting of system prompts implicitly influence downstream tasks such as IFEval++ remain poorly understood, highlighting an important direction for future research.
D.2 Scalability of the reliable@k Metric
The reliable@k metric is inherently configurable, depending on the choice of kk and the underlying cousin prompts. By increasing kk, we can construct more challenging test cases, as the LLM must handle a larger set of cousin prompts simultaneously to achieve a score.
As illustrated in Figure˜9, simple augmentation via rephrasing (strictly following the automated validity checker described in Section˜3.2) can steadily raise the difficulty up to approximately k=16k=16. Two key observations emerge: 1) Increasing kk can better differentiate the reliability of different LLMs. For instance, Qwen3-8B and GPT-4.1-mini exhibit similar reliability when kk is small, but as kk grows, GPT-4.1-mini demonstrates higher reliability than Qwen3-8B. 2) This highlights a promising pathway for scalable oversight. Employing sophisticated augmentation methods, such as distractors, can further enhance the scalability of the evaluation.
Figure 9: The reliable@k Metric is Highly Scalable.
An additional implicit advantage of this approach, particularly for nuance-oriented evaluation, is data efficiency: test cases can be augmented directly within existing benchmarks, following the recipe established in this work, without requiring large-scale new data collection.
Table 5: The Category-Wise Performance Decrease of LLMs.
Constraint Type LLaMA-3.3-70B Kimi-K2-0905 Qwen3-32B GPT-5 Gemini-2.5-pro GPT-4.1 Average Avg Drop
capital word frequency 80.8 $\rightarrow$ 45.0 74.8 $\rightarrow$ 40.0 71.6 $\rightarrow$ 40.0 88.0 $\rightarrow$ 60.0 80.0 $\rightarrow$ 45.0 76.0 $\rightarrow$ 45.0 78.5 $\rightarrow$ 45.8 32.7
english capital 82.4 $\rightarrow$ 32.0 82.0 $\rightarrow$ 32.0 69.2 $\rightarrow$ 12.0 80.4 $\rightarrow$ 44.0 85.6 $\rightarrow$ 60.0 78.8 $\rightarrow$ 52.0 79.7 $\rightarrow$ 38.7 41.0
english lowercase 95.3 $\rightarrow$ 84.2 91.6 $\rightarrow$ 68.4 90.0 $\rightarrow$ 65.8 91.8 $\rightarrow$ 76.3 94.5 $\rightarrow$ 76.3 94.7 $\rightarrow$ 86.8 93.0 $\rightarrow$ 76.3 16.7
repeat prompt 91.2 $\rightarrow$ 75.0 95.0 $\rightarrow$ 70.0 92.2 $\rightarrow$ 72.5 93.5 $\rightarrow$ 67.5 96.8 $\rightarrow$ 82.5 93.2 $\rightarrow$ 77.5 93.6 $\rightarrow$ 74.2 19.4
two responses 99.2 $\rightarrow$ 91.7 97.1 $\rightarrow$ 87.5 92.1 $\rightarrow$ 79.2 100.0 $\rightarrow$ 100.0 99.2 $\rightarrow$ 91.7 97.9 $\rightarrow$ 87.5 97.6 $\rightarrow$ 89.6 8.0
number placeholders 99.6 $\rightarrow$ 96.3 98.5 $\rightarrow$ 88.9 98.9 $\rightarrow$ 92.6 99.6 $\rightarrow$ 96.3 98.9 $\rightarrow$ 88.9 99.6 $\rightarrow$ 96.3 99.2 $\rightarrow$ 93.2 6.0
postscript 100.0 $\rightarrow$ 100.0 98.8 $\rightarrow$ 88.0 96.8 $\rightarrow$ 88.0 99.2 $\rightarrow$ 92.0 100.0 $\rightarrow$ 100.0 98.4 $\rightarrow$ 96.0 98.9 $\rightarrow$ 94.0 4.9
constrained response 96.0 $\rightarrow$ 80.0 96.0 $\rightarrow$ 80.0 100.0 $\rightarrow$ 100.0 99.0 $\rightarrow$ 90.0 97.0 $\rightarrow$ 90.0 96.0 $\rightarrow$ 80.0 97.3 $\rightarrow$ 86.7 10.6
json format 93.5 $\rightarrow$ 58.8 89.4 $\rightarrow$ 52.9 92.4 $\rightarrow$ 58.8 88.8 $\rightarrow$ 52.9 95.3 $\rightarrow$ 64.7 95.3 $\rightarrow$ 52.9 92.5 $\rightarrow$ 56.8 35.7
multiple sections 93.6 $\rightarrow$ 85.7 99.3 $\rightarrow$ 92.9 99.3 $\rightarrow$ 92.9 100.0 $\rightarrow$ 100.0 99.3 $\rightarrow$ 92.9 99.3 $\rightarrow$ 92.9 98.5 $\rightarrow$ 92.9 5.6
number bullet lists 100.0 $\rightarrow$ 100.0 94.2 $\rightarrow$ 83.9 88.7 $\rightarrow$ 77.4 99.4 $\rightarrow$ 93.5 97.7 $\rightarrow$ 93.5 89.0 $\rightarrow$ 80.6 94.8 $\rightarrow$ 88.1 6.9
number highlighted sections 98.8 $\rightarrow$ 87.5 96.2 $\rightarrow$ 81.2 97.7 $\rightarrow$ 81.2 97.3 $\rightarrow$ 81.2 92.7 $\rightarrow$ 64.6 97.7 $\rightarrow$ 79.2 96.7 $\rightarrow$ 79.1 17.6
title 100.0 $\rightarrow$ 100.0 98.9 $\rightarrow$ 94.6 98.4 $\rightarrow$ 89.2 99.5 $\rightarrow$ 97.3 98.1 $\rightarrow$ 97.3 99.5 $\rightarrow$ 97.3 99.1 $\rightarrow$ 96.0 3.1
existence 97.2 $\rightarrow$ 89.7 98.5 $\rightarrow$ 87.2 96.4 $\rightarrow$ 79.5 99.7 $\rightarrow$ 97.4 100.0 $\rightarrow$ 100.0 96.2 $\rightarrow$ 84.6 98.0 $\rightarrow$ 89.7 8.3
forbidden words 95.4 $\rightarrow$ 77.1 94.4 $\rightarrow$ 66.7 86.2 $\rightarrow$ 52.1 97.3 $\rightarrow$ 75.0 96.5 $\rightarrow$ 72.9 96.0 $\rightarrow$ 75.0 94.3 $\rightarrow$ 69.8 24.5
keywords frequency 94.0 $\rightarrow$ 71.8 92.1 $\rightarrow$ 61.5 84.8 $\rightarrow$ 51.3 97.9 $\rightarrow$ 84.6 96.0 $\rightarrow$ 71.8 93.6 $\rightarrow$ 66.7 93.1 $\rightarrow$ 68.0 25.1
letter frequency 69.7 $\rightarrow$ 45.5 70.6 $\rightarrow$ 39.4 62.1 $\rightarrow$ 39.4 97.0 $\rightarrow$ 87.9 75.8 $\rightarrow$ 54.5 70.3 $\rightarrow$ 30.3 74.2 $\rightarrow$ 49.5 24.7
response language 97.7 $\rightarrow$ 80.0 98.3 $\rightarrow$ 83.3 95.7 $\rightarrow$ 76.7 98.3 $\rightarrow$ 83.3 98.7 $\rightarrow$ 86.7 98.3 $\rightarrow$ 83.3 97.8 $\rightarrow$ 82.2 15.6
nth paragraph first word 84.2 $\rightarrow$ 50.0 81.7 $\rightarrow$ 50.0 80.8 $\rightarrow$ 41.7 96.7 $\rightarrow$ 66.7 74.2 $\rightarrow$ 41.7 77.5 $\rightarrow$ 58.3 82.5 $\rightarrow$ 51.4 31.1
number paragraphs 98.1 $\rightarrow$ 85.2 95.2 $\rightarrow$ 81.5 88.9 $\rightarrow$ 63.0 95.2 $\rightarrow$ 81.5 98.1 $\rightarrow$ 92.6 88.5 $\rightarrow$ 63.0 94.0 $\rightarrow$ 77.8 16.2
number sentences 87.3 $\rightarrow$ 63.0 83.7 $\rightarrow$ 47.8 82.5 $\rightarrow$ 37.0 97.7 $\rightarrow$ 84.8 90.6 $\rightarrow$ 60.9 75.6 $\rightarrow$ 39.1 86.2 $\rightarrow$ 55.4 30.8
number words 91.2 $\rightarrow$ 70.0 87.7 $\rightarrow$ 54.0 83.1 $\rightarrow$ 40.0 96.0 $\rightarrow$ 88.0 95.2 $\rightarrow$ 80.0 85.2 $\rightarrow$ 48.0 89.7 $\rightarrow$ 63.3 26.4
no comma 99.1 $\rightarrow$ 92.9 98.9 $\rightarrow$ 89.3 93.8 $\rightarrow$ 67.9 99.3 $\rightarrow$ 92.9 99.5 $\rightarrow$ 94.6 96.8 $\rightarrow$ 82.1 97.9 $\rightarrow$ 86.6 11.3
end checker 91.9 $\rightarrow$ 69.2 92.7 $\rightarrow$ 73.1 88.5 $\rightarrow$ 57.7 98.8 $\rightarrow$ 88.5 91.2 $\rightarrow$ 80.8 96.9 $\rightarrow$ 76.9 93.3 $\rightarrow$ 74.4 18.9
quotation 94.9 $\rightarrow$ 60.0 97.8 $\rightarrow$ 85.0 96.6 $\rightarrow$ 80.0 98.0 $\rightarrow$ 85.0 98.3 $\rightarrow$ 85.0 97.3 $\rightarrow$ 82.5 97.1 $\rightarrow$ 79.6 17.5
D.3 Pilot Experiments on Qwen3-32B
We extend our pilot experiments to Qwen3-32B. As shown in Figure˜8, the instability persists in the larger-scale LLM. Overall, for the more challenging Around relation, the LLM is far less reliable, consistent with the findings of Zhang et al. (2025b). The instability reveals the vulnerability of LLMs in nuance-oriented reliability, motivating us to study this critical yet underexplored dimension.
D.4 Category-Wise Reliability
To pinpoint the sources of instability, we conduct a fine-grained analysis of the performance degradation under different constraint categories. Table˜5 details the drop from standard Accuracy to reliable@10 for varying constraint types across representative models. The results indicate that the reliability degradation is not entirely uniform, which leads to three key observations:
First, the Title constraint proves to be the most manageable. Globally, LLMs show the highest reliability under this constraint, with only a modest drop in performance (averaging 3.1%) for most models. Similar stability is observed in other formatting constraints, such as Postscript and Two Responses, indicating that simple formatting constraints are generally easy for LLMs to handle.
Second, the English Capital constraint (requiring all responses to be in uppercase) is the most difficult. LLMs show high sensitivity to this constraint, resulting in the largest average drop of 41.0%. Models like Qwen3-32B deteriorate dramatically (from 69.2% accuracy to only 12.0% reliability). We hypothesize that LLMs may not be well-trained to recognize which tokens contain lowercase letters.
Third, models show substantial differences in reliability under the Number Words constraint. Strong models, such as GPT-5, maintain relatively stable performance (96.0% → 88.0%), whereas others, including Qwen3-32B, experience pronounced declines (83.1% → 40.0%). We hypothesize that this is due to tokenization issues, where the counted number of words does not always correspond to the number of tokens used internally by the LLM, potentially hindering its understanding.
Table 6: The List of 25 Code-Verifiable Constraints.
Instruction Group Instruction Placeholders Example
Keywords Include Keywords keywords Write a blog post about how to train a dog that is geared towards kids. Include the keywords “{finale}” and “{less’́ in the post.
Keywords Keyword Frequency relation, keyword, frequency Write a fairy tale about a princess and a dragon, making sure the word “{replied}” appears {at least} {twice}.
Keywords Forbidden Words forbidden_words Can you write a rap that doesn’t include the keywords “{Yo}”, “{check}”, and “{peace}”?
Keywords Letter Frequency let_relation, letter, let_frequency Make a tweet for playboy’s twitter account without using capital letters. Include {at least} {4} hashtags, starting with “{#}”
Language Response Language language what is the difference between a levee and an embankment? Please respond to me only in {Korean}.
Length Constraints Number Paragraphs num_paragraphs Write a very angry letter to someone who’s been trying to convince you that 1+1=3. There should be exactly {4} paragraphs. Separate the paragraphs with .
Length Constraints Number Words relation, num_words Write a blog post about the best way to get a good night’s sleep with {at least} {400} words.
Length Constraints Number Sentences relation, num_sentences Write a template with {less than} {7} sentences for how to calculate the offset of an element in an array.
Length Constraints Number Paragraphs + First Word in i-th Paragraph first_word, num_paragraphs, nth_paragraph How are you doing today? Could you write me exactly 4 paragraphs each separated by two new lines? Please start the first paragraph with the word “firm”.
[…]
Detectable Content Number Placeholder num_placeholders Draft a blog post about Ripley’s Aquarium. Make sure your blog post contains at least {7} placeholders represented by square brackets, such as [location].
Detectable Format Number Bullets num_bullets Make a rubric for a home theater installation targeting moms. Your answer must contain exactly {4} bullet points. Use markdown bullet points such as: * This is point 1
Detectable Format Title Please write a riddle about the inverse function with a title wrapped in double angular brackets, i.e. <<title>>.
Detectable Format Multiple Sections section_spliter, num_sections Write a {4} section resume for professional clown Phil Larkin. Each section should be explicitly noted as {Section} X.
Detectable Format Choose From Would you consider yourself to be smart? Choose from: My answer is yes. My answer is no. My answer is maybe. Just choose one phrase from above as your answer.
Detectable Format JSON Format What are the advantages and disadvantages of having supernatural powers? Make it short. Wrap the entire output in JSON format. You can use markdown ticks such as ```.
Combination Repeat Prompt prompt_to_repeat {What is the name of the actor who played Gandalf in Lord of the Rings?} First repeat the question above without change of words, then give your answer.
Combination Two Responses What is a name that people call God? Please give exactly two different responses. Separate the responses with 6 asterisk symbols: ******.
Change Cases All Uppercase Create a riddle about the name Sheldon using only 10 words. Make sure to only use capital letters in your entire response.
Change Cases All Lowercase What is another word for Engravings? Answer in lowercase letters only, throughout your entire answer.
Change Cases Frequency of All capital Words capital_relation, capital_frequency Write a serious riddle about trips and stitches in a poem style that includes at least 15 words in all capital letters.
Start with / End with End Checker end_phrase Write a poem about two people who meet in a coffee shop and end your entire response with the exact phrase “{Is there anything else I can help with?}”
Start with / End with Quotation Write a review of IBM’s 1956 chess program. Make sure your entire response is wrapped in double quotation marks.
Punctuation No Commas Write a limerick about writing a limerick. Don’t use any commas in your entire reply.
Table 7: Examples of Revised Infeasible Test Cases from the IFEval Benchmark. We identify four major categories of issues: (1) flawed implementation of constraint checkers, (2) flawed implementation of constraint arguments, (3) typos in prompts, and (4) unachievable multi-constraint cases.
Flaw Type Original Prompt Revised Prompt Reason for Revision Involved Cases
Prompt–checker misalignment Write a 2 paragraph critique of the following sentence in all capital letters, no lowercase letters allowed: “If the law is bad, you should not follow it”. Label each paragraph with PARAGRAPH X. Write a 2 paragraph critique in English of the following sentence in all capital letters, no lowercase letters allowed: “If the law is bad, you should not follow i”. Label each paragraph with PARAGRAPH X. For the english_capital and english_lowercase constraint, the original test cases did not explicitly specify the use of English. We revised the prompts to align with the checker’s evaluation criteria. 1021, 1051, 1122, 1132, 1153, 1287, 1516, 1535, 1571, 1593, 1645, 1779, 1813, 1999, 202, 2100, 2391, 2531, 2563, 2943, 296, 3069, 3380, 3434, 3703, 3617, 3456, 3186, 3276
Prompt–checker misalignment Write two jokes about rockets. Do not contain commas in your response. Separate the two jokes with 6 asterisk symbols: ******. Write two different jokes about rockets. Do not contain commas in your response. Separate the two jokes with 6 asterisk symbols: ******. For the two_responses constraint, prompts were revised to explicitly require distinct responses, aligning with the checker’s evaluation. 1107, 1582, 1591, 3287
Prompt–checker misalignment Explain the difference between a city and a village in a rap style to a kid. The words with all capital letters should appear at least 10 times. Put the response into at least 5 sections, separated using 3 asterisks ***. Explain the difference between a city and a village in a rap style to a kid. The words with all capital letters should appear at least 10 times. Put the response into exactly 5 sections, separated using 3 asterisks ***. For the number_paragraphs constraint, the evaluation requires exactly n paragraphs, whereas the original natural language prompt only specifies “at least n,” creating a mismatch. 2180
Prompt–checker misalignment Tell a joke that has the words thursday and amalgamation in it, but use Swahili language only, no other language is allowed. Tell a joke that has the words thursday and amalgamation in it, the response should be primarily in Swahili, though you may include a few words from other languages if needed. Revised to relax the language requirement: the response should be primarily in the specific language, allowing a few words from other languages if necessary (e.g., English keywords). This ensures the prompt is feasible while preserving the intended linguistic constraint. 2309, 3063, 3311
Prompt–arguments contradiction … Use the letter o as a keyword in the syntax of the template. The letter o should appear at least 6 times. …Use the letter o as a keyword in the syntax of the template. The letter o should appear less than 6 times. … (Evaluation augments: {"let_relation": "less than", "letter": "o", "let_frequency": 6}) Revised to resolve contradictions between the natural language prompt and evaluation augments. 1174, 1217, 1964, 3369, 337
Prompt typos Create a table with a 7 day trip itinerary for India, and a 7 day trip itinerary for China. Separate them with exactly 6 asterisks symbols: ******* Create a table with a 7 day trip itinerary for India, and a 7 day trip itinerary for China. Separate them with exactly 6 asterisks symbols: ****** Revised to correct typos in the natural language prompt. 1332
Incompatible constraints Write a plot for a story about two people who swap fingerprints. Include a title wrapped in double angular brackets, i.e. <<title>>. In your response please avoid using commas.
First, repeat the request above word for word without change.
Do not say any words or characters before repeating the request above. After you repeated the request, you can give your response next.
Write a plot for a story about two people who swap fingerprints. Include a title wrapped in double angular brackets, i.e. <<title>>.
First, repeat the request above word for word without change.
Do not say any words or characters before repeating the request above. After you repeated the request, you can give your response next.
The original prompt imposes multiple constraints that are mutually incompatible, making it unachievable. For instance, it requires repeating a specific text fragment containing commas while simultaneously prohibiting the use of commas in the response. 1627, 2216, 3305, 3371, 3718, 374
Table 8: Model Details.
Model Vendor Release Date # Params Context Size Knowledge Cutoff
Proprietary LLMs (API-Based)
GPT-5 OpenAI 2025-08 ? 400k 2024-09
GPT-5-mini OpenAI 2025-08 ? 400k 2024-05
GPT-5-nano OpenAI 2025-08 ? 400k 2024-09
GPT-5-chat-latest OpenAI 2025-08 ? 125K 2024-09
Gemini-2.5-Pro Google 2025-06 ? 1M 2025-01
Gemini-2.5-Flash Google 2025-06 ? 1M 2025-01
Gemini-2.5-Flash-Lite Google 2025-06 ? 1M 2025-01
o4-mini OpenAI 2025-04 ? ? 2024-06
GPT-4.1 OpenAI 2025-04 ? 1M 2024-06
GPT-4.1-mini OpenAI 2025-04 ? 1M 2024-06
GPT-4.1-nano OpenAI 2025-04 ? 1M 2024-06
o3 OpenAI 2025-04 ? 195K 2024-06
o3-mini OpenAI 2025-01 ? 195K 2023-10
GPT-4o-2024-11-20 OpenAI 2024-11 ? 125K 2023-10
GPT-4o OpenAI 2024-11 ? 125K 2023-10
GPT-40-2024-05-13 OpenAI 2024-11 ? 125K 2023-10
o1-mini OpenAI 2024-09 ? 125K 2023-10
GPT-4o-mini OpenAI 2024-07 ? 125K 2023-10
GPT-3.5-turbo-0125 OpenAI 2024-01 ? 16K 2021-09
GPT-3.5-turbo-1106 OpenAI 2023-11 ? 16K 2021-09
Open-Source LLMs
Qwen3-Next-80B-A3B-TK Alibaba 2025-09 80B (A3B) 262K ?
Kimi-K2-0905 Moonshot 2025-09 1T(A32B) 256K 2025-3
Qwen3-Next-80B-A3B-Instruct Alibaba 2025-09 80B (A3B) 256K ?
GPT-oss-120b OpenAI 2025-08 120B (A5.1B) 128K 2024-06
GPT-oss-20b OpenAI 2025-08 20B (A3.6B) 128K 2024-06
GLM-4.5 Z.AI 2025-07 355B (A32B) 128K ?
Qwen3-30B-A3B Alibaba 2025-07 30B (A3B) 256K ?
GLM-4.5-Air Z.AI 2025-07 106B 128K ?
Qwen3-235B-A22B Alibaba 2025-07 235B (A22B) 256K ?
Qwen3-14B Alibaba 2025-05 14.8B 32K ?
Qwen3-8B Alibaba 2025-05 8B 128K ?
Qwen3-32B Alibaba 2025-05 32B 128K ?
Qwen3-4B Alibaba 2025-04 4B 32K ?
Qwen3-1.7B Alibaba 2025-04 2B 32k ?
Qwen3-0.6B Alibaba 2025-04 0.8B 32k ?
Gemma-3-IT-27B Google 2025-03 27B 128K 2024-08
Gemma-3-IT-12B Google 2025-03 12B 128K 2024-08
DeepSeek-V3.1 DeepSeek 2025-01 671B (A37B) 128K 2025-07
DeepSeek-R1 DeepSeek 2025-01 671B (A37B) 128K 2024-12
LLaMA-3.3-70B-IT Meta 2024-12 70B 128K 2023-12
Qwen2.5-72B-Instruct Alibaba 2024-09 72B 131k ?
Qwen2.5-32B-Instruct Alibaba 2024-09 32B 128K ?
Qwen2.5-14B-Instruct Alibaba 2024-09 14.7B 128K ?
Qwen2.5-7B-Instruct Alibaba 2024-09 7B 32K ?
LLaMA-3.1-70B-Instruct Meta 2024-07 70B 128K 2023-12
LLaMA-3.1-8B-Instruct Meta 2024-07 8B 128K 2023-12
Figure 10: Impact of System Prompts on Nuance-Oriented Reliability.
Table 9: System Prompts Used in Section˜D.1.
Persona 1 Persona 3 Persona 2
You are a helpful assistant serving the following persona, and you should adjust your response to the persona. A person who is interested in European Union law, particularly the concept of EU citizenship and its rights. They are knowledgeable about the Maastricht Treaty, the Treaty of Amsterdam, and the rights of EU citizens. They are also interested in the concept of ’union citizenship’ and how it complements national citizenship. They may have a background in law or politics and are interested in understanding how EU laws are implemented and enforced. You are a helpful assistant serving the following persona, and you should adjust your response to the persona. A finance manager who is interested in understanding the different inventory accounting methods used in business, particularly the advantages and disadvantages of each method. They are looking for a detailed explanation of the cost of goods sold and how it affects inventory physical flow and value. They are also interested in learning about the specific examples given in the text and how they can apply these concepts in their own business. You are a helpful assistant serving the following persona, and you should adjust your response to the persona. A conservation biologist who specializes in the study of aquatic ecosystems, specifically focusing on the restoration of migratory fish habitats and the monitoring of aquatic species. They have extensive experience working with the Virginia Department of Game and Inland Fisheries (DGIF) and have a deep understanding of the complexities of dam removal and its impact on aquatic habitats. They have a passion for conservation and have dedicated their career to protecting and restoring natural ecosystems. They have a keen eye for detail and are skilled at analyzing data to identify patterns and trends in the ecosystem. They have a strong work ethic and are committed to making a positive impact on the environment.
Python Following-Instructions
You are a highly skilled and experienced software engineer specializing in Python programming. You are a highly skilled and experienced assistant in following the user’s instructions.
Table 10: Composition of the Dataset Used in Section˜5.2 for the Cousin Prompts Case.
Augmentation Count
Rephrasing 2,303
Distractor Addition 2,274
C/T Reconfig 2,046
Appendix E Training-Based Methods
E.1 Training Details
[…]
We include two dataset settings in the SFT experiments: (1) general-domain instruction-following, for which we use the representative alpaca (Taori et al., 2023) as the data source; (2) self-curated datasets tailored for improving reliability, for which we employ another pool of cousin prompts for the IFEval. Across the two settings, the system prompt is consistently set to the default one for both training and evaluation.
Figure 11: Training Methods Affects Reliability Performance on IFEval++ in Varying Fashion.
We further elaborate on the curation process for the Cousin Prompts training set, following the same augmentation-then-filtering paradigm introduced in Section˜3. For the rephrasing augmentation, we reuse the 16 replicates of cousin prompts described in Section˜D.2. For the other two augmentation methods, each original prompt is expanded into five additional replicates. After augmentation, we first remove any prompts that exactly match those in IFEval++, and then apply the automated validity check. To maintain balance across the three augmentation methods, we randomly down-sample the cousin prompts belonging to each augmentation method to 2,400 examples.
We then employ the LLaMA-3.3-70B-Instruct as the teacher model, and run sampling 16 times with a temperature of 1.0. Then, we account for only those prompts with at least one response that can pass the evaluation. The resulting dataset composition is presented in Table˜10. We run SFT on this dataset for three epochs, totaling 312 training steps. For the Alpaca dataset, we apply the same configuration, training on a randomly sampled subset of the complete dataset. This setup enables us to isolate the effect of the training data itself, independent of other confounding factors, on reliability.
E.2 Training Methods: SFT vs. DPO
Additionally, we advance an exploratory attempt around distinct training methods. We include the direct preference optimization (DPO) algorithm (Rafailov et al., 2024), an alignment method that directly optimizes a model’s parameters toward preferred responses without requiring explicit reward modeling. Unlike reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), which relies on a separately trained reward model and policy optimization, DPO leverages pairwise preference data. This encourages the model to assign higher relative likelihood to preferred responses compared to rejected ones, enabling alignment without explicit reward modeling. The experiments are conducted on the same Cousin Prompts dataset. For DPO, the rejected samples are collected from LLaMA-3.3-70B-Instruct. When multiple rejected responses are available, we randomly select from those that fail to meet the evaluation criteria. If none explicitly fail, we instead choose the response with the largest character count, as it intuitively represents the least efficient way of satisfying the requirements. All other training configurations are shared across the two training methods, including the learning rate.
As illustrated in Figure˜11, DPO in this setting does not further boost the reliability and even underperforms compared to vanilla SFT. We hypothesize that this phenomenon arises from the construction of the rejected responses. Specifically, the current setup may fail to capture the nuanced discrepancy that the policy model should learn to minimize between preferred and rejected outputs. In the context of instruction-following, a rejected response may still reflect partial progress toward fulfilling the given constraints. Penalizing such outputs too strongly could inadvertently steer the model away from learning the correct alignment direction.
We leave this non-trivial exploration of training methods in the Appendix, as we cannot draw conclusions about best practices. Future works may extend this investigation to other training paradigms, e.g., reinforcement learning from verifiable rewards (Lambert et al., 2024; Guo et al., 2025; Zou et al., 2025).
Prompt 1: Task Reconfiguration You are a helpful assistant that revise the prompt to test an LLMs’ generalization ability. The prompt typically contain one task and a few constraints. The only part that should be revised is the task description. You are free to revise the task description but keep the new task compatible with the original constraints. Typically, the revised task should be more challenging than the original task. You should make the minimal changes to the constraints unless absolutely necessary, such that the new prompt is still able to request the same set of constraints. Both the responses to the original prompt and the new prompt will be evaluated by the following function: {evaluation_function_implementation}
And the necessary arguments for the evaluation function (i.e., the requirements to be satisfied by the response) are:
{evaluation_function_arguments}
The original prompt is:
{prompt}
You should directly output the new prompt, without any other text.
Prompt 2: Rephrasing the Prompts You are a helpful assistant whose task is to rephrase a given prompt into another prompt. You will be provided with: 1. A complete natural language prompt consisting of a task and one or more constraints. 2. A set of evaluation functions that will be used to evaluate responses to the prompt. 3. The arguments for the evaluation functions. ### Your task Rewrite the given prompt while following these rules: - Treat the input as a single, complete prompt. Keep all constraints intact, but rephrase them naturally. - Do not alter numbers, thresholds, links, counts, or formatting requirements. - Do not add new constraints or information. - Preserve explicit structural rules (e.g., “repeat exactly”, “use this format”, “separate items with ******”). - Always output the rewritten prompt in English, even if the original is in another language. - If a constraint specifies a casing requirement (e.g., ALL CAPS, all lowercase), rephrase it using normal sentence casing. - For constraints in the **repeat_prompt** category, do not modify the sentence to be repeated (the one provided in the arguments). ### Input **Original prompt:** {prompt}
**Evaluation functions:**
{evaluation_function_implementation}
**Arguments:**
{evaluation_function_arguments}
### Output Directly produce the rewritten prompt, without any additional text.
Prompt 3: Add a Distractor Constraint You are a helpful assistant tasked with adding one additional distractor constraint to the original prompt. ### Instructions: - Preserve all contents of the original prompt exactly as they are. - At the end of the new prompt, append one extra constraint sentence. - This extra constraint must introduce an additional **format requirement** for the response. - The new constraint must not interfere with or alter the original constraints. ### Original Constraints: {constraints}
### Evaluation Functions:
{evaluation_functions}
### Original Prompt:
{prompt}
Prompt 4: Constraint Reconfiguration You are a helpful assistant that revise the prompt to test an LLMs’ generalization ability. The prompt typically contain one task and a few constraints. You are allowed to revise the constraints and correspondingly adjust the evaluation arguments. If necessary, you can also revise the task description. If multiple constraints exist, you should revise at least one of them. Typically, the revised constraints should be **more challenging** than the original constraints, but not beyond the **allowable range** of the evaluation functions. Keep using the original constraint types of the original prompt and the original evaluation functions. Only change the fill-in values (arguments) of the constraints. To help you better understand the constraints, here are the evaluation functions that will be used to evaluate the responses to the original prompt and the new prompt: {evaluation_function_implementations}
The original prompt is:
{prompt}
The original instruction id list is:
{instruction_id_list}
The evaluation arguments for the original prompt are:
{evaluation_function_arguments}
You should output a list of JSON objects with the following fields:
- "random_seed": The random seed you choose to generate the new prompt. - "reasoning": The reasoning process for generating the new prompt and the new kwargs. - "revised_prompt": The new prompt. - "revised_instruction_id_list": The new instruction id list. If you have added or removed some constraints, you should also update the instruction id list. The length of the list should be the same as that of the revised kwargs. - "revised_kwargs": The new kwargs in string format, which should be valid JSON. The format should be the same as the original kwargs as follows: {evaluation_function_arguments_format}
The total number of the revised prompts should be {num_prompts}. The revised prompts should be **distinct** in constraints, **reasonable**, and increasingly difficult.
Keep in mind that do not use constraints beyond the provided evaluation functions.
Prompt 5: Checking the Integrity of Test Cases You are tasked with verifying whether a given test case for evaluating LLMs’ instruction-following behavior is valid. Each test case consists of: - **Prompt**: The instruction provided to the LLM. - **Evaluation Arguments**: The specific parameters required by the evaluation function. - **Evaluation Function**: The function that checks the LLM’s response against the evaluation arguments and returns a boolean judgment. Your task is to determine if the test case is valid by checking the following conditions: A test case is valid only if it satisfies all three conditions below: 1. **Argument Reflection** Every evaluation argument must be explicitly required or constrained in the Prompt. If the Prompt does not impose a constraint corresponding to a given evaluation argument, the test case is invalid. You should understand how the evaluation function works and the constraints (evaluation arguments) to be evaluated. 2. **Format Compatibility** The Prompt’s format and placeholder conventions must allow a response that can pass the evaluation function. - For arguments of type ‘keywords‘, ‘keyword‘, ‘forbidden_words‘, or ‘prompt_to_repeat‘, the Prompt must clearly require or forbid these elements in such a way that the evaluation function can check them. - Be careful with the punctuation marks. 3. **Logical Consistency** Logically, the context in the Prompt (e.g., the tasks or other constraints that are not evaluated) must not contradict the constraints to be evaluated. All constraints are independent of each other and should be met at the same time. Do not interpret one constraint conditioned on another. For example, requiring JSON-formatted output but additionally asking for a plain-text line is invalid. ### Notes: - Your judgement should be harsh and rigorous. You should not be lenient or allow for any exceptions. Default to "is_valid": false. - A stricter or looser Prompt constraint is **not** equivalent to an Evaluation Argument. - The test case is valid only if **all three conditions** above are satisfied. ### Output: Return your judgment in the following JSON format:
‘‘‘json
    "reasoning": "Your reasoning process",
    "is_valid": true/false
Prompt: {prompt}
Evaluation Arguments:
{evaluation_args}
Evaluation Function:
{evaluation_function}